The speed of a boat in still water is 15 km/hr. It needs four more hours to travel 63 km against the current of a river than it needs to travel down the river. Determine the speed of the current of the river.


1
Expert's answer
2021-04-15T07:20:05-0400

Let "x" is the speed of the current of the river.

Time against the current minus time down the river is 4.

Equation:


"\frac{63}{15-x}-\frac{63}{15+x}=4"


Then we multiply equation by "15^2-x^2=(15-x)(15+x)" :


"63(15+x)-63(15-x)=4(15^2-x^2)"


"4x^2+126x-4*225=0"


"2x^2+63x-450=0"


"x=\frac{-63+\sqrt{63^2+4*2*450}}{2}="


"=\frac{-63+\sqrt{3969+3600}}{4}="


"=\frac{-63+87}{4}=6"


Answer: "x=6"